OpenStudy (anonymous):

Identify the minimum y=x^2-4x-32

OpenStudy (anonymous):

y=x^2-4x-32 if you solve for x you get 8 and -4 the max should be in the middle of those 2 points right?

OpenStudy (anonymous):

2

sam (.sam.):

or just differentiate, set dy/dx=0 y=x^2-4x-32 dy/dx=2x-4=0 x=2 then find y

OpenStudy (anonymous):

assuming he knows calculus

OpenStudy (anonymous):

but yes that is easier

sam (.sam.):

we could use completing the square x^2-4x-32 (x-2)^2-4-32 (x-2)^2-36 symmetry at x=2

OpenStudy (anonymous):

Yup he should definately know CTS